OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 2 - 12 + 72 - 432 + ...

OpenStudy (anonymous):

\[\sum_{x=0}^{\infty}\left( 2 \right)\left( -6 \right)^{x}\]

OpenStudy (anonymous):

All good now, @tholn ?

OpenStudy (anonymous):

thx man!!!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @tholn

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

It's becaue (-6)^x will alternate signs. @Loser66

OpenStudy (anonymous):

And it will start at 1, then -6, then 36, then -216. And each is multiplied by 2. @Loser66

OpenStudy (anonymous):

It's too fundamental. the "-" in the ( ) will provide alternation.

OpenStudy (anonymous):

The whole series, each term is multiplied by 2, so the "2" really could have gone outside the summation notation. np.

OpenStudy (anonymous):

One last thing. The power of 6 or -6 is seen by dividing a term by the previous one. That's the real key.

OpenStudy (loser66):

ok, flexibility, too.

OpenStudy (anonymous):

So with those 3 things, and that's all there really is: 1) each term multiplied by 2, 2) power of 6, and 3) alternating. That's all there is to it.